Sales of Blistre Autos are 380,000, variable cost is 200,000, fixed cost is 80,000 tax rate is 40%. Calculate the operating leverage of the company.

1.11 times

0.80 times

1.80 times

3.00 times
Answer #1

Contribution margin = Sales - Variable costs

= 380,000 - 200,000 =180,000

Operating income = Contribution margin - Fixed cost

= 180,000 - 80,000 = 100,000

Degree of operating leverage = Contribution margin /Operating income

= 180,000/100,000

= 1.8 times

Option C is the answer